OpenStudy (anonymous):

In 5 years, Dad will be three times as old as his daughter Jill will be then. If the sum of their present ages is 50, how old are they now? If x + 5 is Jill's age five years from now, which of the following equations could be used to solve the problem? 3x + 5 = 55 - x 3x + 15 = 55 - x 3x + 15 = 45 - x

OpenStudy (solomonzelman):

x+y=50 (x+5)3=y=5 ---> 3x+15=y+5 ---> 3x+10=y substitute, x+(3x+10)=50 4x+10=50 4x=40 x=10 this is the way I would do it.

OpenStudy (solomonzelman):

I think it is B.

OpenStudy (anonymous):

thank you

OpenStudy (solomonzelman):

Anytime!

OpenStudy (anonymous):

what about this one? The length of a rectangle is 2 more than twice the width. Find the dimensions of the rectangle if the perimeter is 76. If x represents the width of the rectangle, then which expression represents the perimeter? 2x + 2 4x + 4 6x + 4

OpenStudy (solomonzelman):

l-2=2w lw=76 can you solve this system?

OpenStudy (solomonzelman):

l -- lenght w -- width

OpenStudy (anonymous):

sorry no

OpenStudy (solomonzelman):

l-2=2w ----> l=2w + 2 lw=76 substitute, w(2w + 2)=76 2w^2+2w-76=0 use Completing the square, factoring, or quadratic formula to solve this. can you?

OpenStudy (solomonzelman):

2w^2+2w-76=0 divide the entire equation by 2 w^2+w-32=0 can you solve this?